# Connecting Neural Models Latent Geometries with Relative Geodesic Representations

### v1

@@ 5.2 Zero-shot stitching of vision foundation models
Rel(Cosine) Moschella et al. (2023) $0.907\pm 0.09$ $0.775\pm 0.132$ $\mathbf{0.549}\pm 0.152$ $0.531\pm 0.188$ $0.384\pm 0.115$
RelGeo(Pullback) $\mathbf{0.955}\pm 0.03$ $\mathbf{0.874}\pm 0.055$ $0.501\pm 0.159$ $\mathbf{0.595}\pm 0.163$ $\mathbf{0.59}\pm 0.054$
@@
Model stitching was introduced in Lenc and Vedaldi (2015) to analyze neural network representational spaces, by training a linear layer to connect different layers and evaluating performance. Here we sidestep the need for trainable stitching layers and consider the zero-shot model stitching task defined in Moschella et al. (2023) to effectively test how components of vision foundation models can be reused. To do this, we leverage the space of relative geodesic representations as a shared compatible space. For the $i$th model $E_{i}$, we train one decoder $D_{i}$ on the relative representations induced by it, then evaluate the performance of using $D_{i}$ to decode the representations of model $E_{j}$, where $E_{j}$ may be a different model. This assesses how much two representation spaces can be merged with respect to the task defined by decoder $D$, e.g., a classification head.
Experimental setting.   We perform experiments on pretrained vision foundation models from Hugging Face, investigating how well we can match representations together with different backbones with classification heads, on the same datasets and models as considered in Section 5.1. Similar to Section 5.1, we compare Rel(Cosine), RelGeo(Pullback) and RelGeo(Diet).
Analysis of results.   The results of the different methods across the different data sets are shown in Table 2, where we average over all possible model pairs. We further show the accuracies of the models on the CUB dataset in Fig. 5. Both RelGeo(Pullback) and RelGeo(Diet) provide strong stitching accuracies, with RelGeo(Pullback) reflecting the benefits of pulling back class specific information. RelGeo(Diet) results still in good accuracies while having very strong MRR metrics 1.
Takeaway: Using geometric relative representations yields good accuracies and good MRRs, avoiding downgrading of performance when performing model stitching while retaining identifiability.
@@ 6 Conclusions and discussion
Limitations and future work.   The accuracy of using the straight line arc length (or energy) approximation can be imprecise in regions of high curvature in the latent space, corresponding to regions further from the training points’ support. Moreover, this could require increasingly smaller step sizes, hurting the efficiency performance of the method. This suggests considering different paths rather than the linear one, and adaptive step sizes, e.g., by estimating the support of the data building KNN graphs in the latent space and forcing the path to not deviate too much from them. By employing the pullback metric from a given output space, the relative geodesics representation has the interesting property of restricting the alignment problem to the information relevant to the decoding task. This could be useful to (i) explore multi-modal alignment Norelli et al. (2023), where it is of interest to capture not only the shared information across modalities, but also the modality-specific information; (ii) to better understand the relation between the representation similarity and decodability Harvey et al. (2024) and the interaction between tasks and learned representations Fumero et al. (2023).
@@ Appendix A Appendix
A.1 Additional details
A.1.1 Mean Reciprocal Rank
@@
Mean Reciprocal Rank (MRR) is a commonly used metric to evaluate the performance of retrieval systems [Moschella et al., 2023]. It measures the effectiveness of a system by calculating the rank of the first relevant item in the search results for each query.
@@
    $\text{MRR}=\frac{1}{Q}\sum_{i=1}^{Q}\frac{1}{r_{i}}$ (6)
@@
Similar to stitching accuracies, MRR is generally asymmetric. However, it can also be made symmetric. Specifically, as MRR is calculated based on a distance matrix $D$, one can make the distance matrix symmetric by setting $D=\frac{1}{2}\left(D^{\top}+D\right)$. In Section 5.1 we reported the symmetric version. Otherwise we report both the original version and the symmetric version, and discriminate between these two by explicitly indicating it when it is symmetric.
A.1.2 Architectural details
We provide here the architectural details of the convolutional Autoencoders employed in experiments in Figures 3 and 4.
@@ A.1.2 Architectural details
Table 5:
A.1.3 RelGeo(Diet) augmentations
@@ A.1.4 Compute resources
The autoencoder stitching and retrieval experiments are performed on an NVIDIA RTX 3080TI GPU.
The experiments concerning vision foundation models are carried out on a compute cluster, with each job using a single NVIDIA A100 GPU and 10 CPU cores and taking several hours.
Preliminary experiments used up more compute resources. It is estimated that we used hundreds of GPU hours.
A.1.5 Geodesic approximation
@@
We trained a lightweight convolutional autoencoder (see Table 6) on both MNIST and CIFAR-10 to obtain the latent representations used in our experiments. For MNIST, the first convolutional layer was adjusted to accept a single input channel; for CIFAR-10 it used three channels. Each model was trained for 30 epochs using the Adam optimizer Kingma and Ba [2017] with a batch size of 64. We set the learning rate to 0.001, and we fixed a random seed of 42 to ensure reproducibility.
Distance computation
After training, we selected 10 samples per class (100 total) in label order from each dataset and encoded them to produce their latent encodings. True geodesics are computed using Stochman library Detlefsen et al. [2021], which wraps the decoder into a pull-back manifold, intializes a spline path between codes, and then optimizes its control points to minimize the Riemannian energy. Geodesic energies are computed as in Eq. 2. Pairwise distances are computed and visualized in Figures 2 and 6, demonstrating the close agreement between the two measures under identical encoding and discretization settings. In Fig. 2, latent dimensions for MNIST and CIFAR are 64 and 128 respectively, while in Fig. 6, latent dimension is 2 for both datasets.
@@
Figure 6: Pairwise latent-space distance matrices for (a) MNIST and (b) CIFAR-10, with latent dimension of 2. In each subfigure, the left heatmap shows the straight-line energy proxy and the right shows the full Riemannian geodesic distances. The Spearman rank correlation between the two measures is 0.99 for MNIST and $\rho=1.00$ for CIFAR-10, demonstrating near-perfect agreement.
(a) MNIST
(b) CIFAR-10
Figure 7: Impact of varying discretization levels on similarity and distance metrics for (a) MNIST and (b) CIFAR-10 datasets. Each subplot shows how Spearman’s $\rho$, Pearson’s $r$, and Euclidean distance change as the number of discretization levels increases.
A.1.6 Autoencoder stitching and retrieval
@@
We trained the lightweight convolutional autoencoder (see Table 6) on MNIST, CIFAR-10, FashionMNIST with 5 different seeds, to obtain the latent representations used in our experiments. For MNIST, and FashionMNIST the first convolutional layer was adjusted to accept a single input channel; for CIFAR-10 it used three channels. Each model was trained for 50 epochs, reaching convergence, using the Adam optimizer Kingma and Ba [2017] with a batch size of 64. We set the learning rate to 0.001.
A.1.7 Vision foundation models
@@
For CIFAR-100, we use coarse labels.
For SVHN, the objective is to predict the cropped digits.
For CUB dataset, we use https://huggingface.co/datasets/birder-project/CUB_200_2011-WDS. The training set is of a small size, and we select $2000$ points as the validation set.
When reporting aggregated MRR metrics in the tables, we always ignore the diagonal numbers as these are generally (close to) $1$.
For all cases where we need to train classification heads, apart from the ones with Diet the heads are trained for $10$ epochs, while the ones with Diet are trained for $50$ epochs. The heads used to obtain the gometric information are trained using learning rate $45e-$ 4 and batch size $64$, while the heads used for stitching was trained using learning rate $41e-$ 4 and batch size $32$.
@@
A.2 Additional results on Vision Foundation models
We provide additional results on vision foundation models. For ablation studies, we focus on the performances of the models on CUB dataset.
A.2.1 Other models
We provide the heatmaps on different datasets in Figure 8, Figure 9, Figure 10 and Figure 11.
Figure 8: CIFAR-10 Accuracies, symmetricized MRR cdist
Figure 9: CIFAR-100 Accuracies, symmetricized MRR cdist
Figure 10: SVHN Accuracies, symmetricized MRR cdist
Figure 11: ImageNet Accuracies, symmetricized MRR cdist
A.2.2 Other evaluation metrics

### v2

@@ 4.2.2 Zero-shot stitching of vision foundation models
Rel(Cosine) (Moschella et al., 2023) $0.907\pm 0.09$ $0.775\pm 0.132$ $\mathbf{0.549}\pm 0.152$ $0.531\pm 0.188$ $0.384\pm 0.115$
RelGeo($L2$) $\mathbf{0.955}\pm 0.03$ $\mathbf{0.874}\pm 0.055$ $0.501\pm 0.159$ $\mathbf{0.595}\pm 0.163$ $\mathbf{0.59}\pm 0.054$
@@
Model stitching was introduced in Lenc and Vedaldi (2015) to analyze neural network representational spaces, by training a linear layer to connect different layers and evaluating performance. Here we sidestep the need for trainable stitching layers and consider the zero-shot model stitching task defined in Moschella et al. (2023) to effectively test how components of vision foundation models can be reused. To do this, we leverage the space of relative geodesic representations as a shared compatible space. For the $i$th model $E_{i}$, we train one decoder $D_{i}$ on the relative representations induced by it, then evaluate the performance of using $D_{i}$ to decode the representations of model $E_{j}$, where $E_{j}$ may be a different model. This assesses how much two representation spaces can be merged with respect to the task defined by the decoder $D$, e.g., a classification head.
Setting.  We perform experiments on pretrained vision foundation models from Hugging Face Transformers (Wolf et al., 2020), investigating how well we can match representations together for classification with different backbones with classification heads, on the same datasets and models as considered in Section 4.2.1, similarly comparing Rel(Cosine), RelGeo($L2$) and RelGeo(Diet).
Analysis of results.  The results of the different methods across the different datasets are shown in Table 2, where we average over all possible model pairs. We further show the accuracies of the models on the CUB dataset in Fig. 5. Both RelGeo($L2$) and RelGeo(Diet) provide strong stitching accuracies, with RelGeo($L2$) reflecting the benefits of pulling back class specific information. RelGeo(Diet) still results in good accuracies while having very strong MRR metrics, as shown in 1.
Takeaway. Relative geodesic representations yield good accuracies and good MRRs, avoiding downgrading of performance when performing model stitching while retaining sample identifiability.
4.2.3 Matching different modalities
In this section we evaluate relative geodesic representations in the multimodal setting.
Figure 6: Matching multimodal models. Symmetricized MRR cosine on Flickr30k. RelGeo(Diet) substantially improves upon Rel(Cosine) in aligning multimodal models.
Setting.  We study the retrieval task in terms of vision foundation models and the text encoders of CLIP (Radford et al., 2021) with both patch 16 and patch 32, using Flickr30k dataset (Young et al., 2014). Keeping the text encoder of CLIP fixed, we swap the vision encoder with the ones of ResNet-50, DINOv2 and ViT, including different patch and model sizes. Due to the lack of class labels, RelGeo($L2$) is not applicable, and we compare RelGeo(Diet) with Rel(Cosine). While we observed that using data augmentations is beneficial for RelGeo(Diet), due to the lack of a principled approach to construct data augmentations on texts corresponding to image augmentations, we do not employ augmentations.
Analysis of results.  The results in terms of symmetricized MRR metric with CLIP with patch 16 are shown in Figure 6. We observe that RelGeo(Diet) yields significantly improved stitching performances upon Rel(Cosine). In Appendix A.4.9 we show the full pairwise matrices of MRR, comprising of the unimodal performances, inter vision models, and text models.
Takeaway. Relative geodesic representations show promising results for obtaining identifiable representations in multimodal scenarios.
@@ 5 Conclusions and discussion
Limitations and future work.  The accuracy of approximating geodesics using straight-line arc length (or energy) can deteriorate in regions of high curvature in the latent space, typically corresponding to areas far from the support of the training data. Moreover, this could require increasingly smaller step sizes, hurting the efficiency performance of the method. This suggests exploring nonlinear paths, and adaptive step sizes, e.g., by estimating the support of the data building KNN graphs in the latent space and forcing the path to not deviate too much from them. By employing the pullback metric from a given output space, the relative geodesic representation has the interesting property of restricting the alignment problem to the information relevant to the decoding task. This could be useful to (i) further explore no training multi-modal alignment (Norelli et al., 2023), where it is of interest to capture not only the shared information across modalities, but also the modality-specific information; (ii) to better understand the relation between the representation similarity and decodability (Harvey et al., 2024) and the interaction between tasks and learned representations (Fumero et al., 2023).
@@ Appendix A Appendix
A.1 Proof of theoretical results
A.1.1 Proof of Proposition 3.1
Proof.
We first prove the first half, i.e. the invariance of Riemannian curve length and energy across reparameterizations of the manifold. This can be proven by observing that the inner product at a point along the curve is invariant across such reparameterizations:
$\bigl\lVert\bm{\dot{x}}\bigr\rVert_{G}$ $=\bm{\dot{x}}^{\top}G\bigl(\bm{x}\bigr)\bm{\dot{x}}=\bigl(\tfrac{d\bm{x}}{d\bm{x^{\prime}}}\bm{\dot{x^{\prime}}}\bigr)^{\top}\bigl(\tfrac{d\bm{x^{\prime}}}{d\bm{x}}\bigr)^{\top}G^{\prime}\bigl(\bm{x^{\prime}}\bigr)\tfrac{d\bm{x^{\prime}}}{d\bm{x}}\tfrac{d\bm{x}}{d\bm{y}}\bm{\dot{x^{\prime}}}$
$=\bm{\dot{x^{\prime}}}^{\top}G^{\prime}\bigl(\bm{x^{\prime}}\bigr)\bm{\dot{x^{\prime}}}=\bigl\lVert\bm{\dot{x^{\prime}}}\bigr\rVert_{G^{\prime}}.$
As such, the length and the energy of the same curves on different manifolds are integrals of the same quantities, hence are equal.
We then prove the second half, i.e. the invariance of Riemannian curve lnegth across reparameterizations of the curve. Based on Equation 4.7 from Hauberg [2025], we have
$L[\bm{\gamma^{\prime}}]$ $=\int_{0}^{1}\bigl\lVert\tfrac{d\bm{\gamma}}{d\tau}\bigr\rVert_{G}\,d\tau=\int_{0}^{1}\bigl\lVert\tfrac{d\bm{\gamma}}{d\tau}\bigr\rVert_{G}\tfrac{\varphi(t)}{t}\,dt$
$=\int_{0}^{1}\bigl\lVert\tfrac{d\bm{\gamma}}{d\tau}\tfrac{\varphi(t)}{t}\bigr\rVert_{G}\,dt=\int_{0}^{1}\bigl\lVert\tfrac{d\bm{\gamma}}{dt}\bigr\rVert_{G}\,dt$
$=L[\bm{\gamma}].$
A.1.2 Proof of Equation 3
Proof.
[…]
A.2 Additional explanations
A.2.1 Details on Diet
Proposed as a self-supervised learning method [Ibrahim et al., 2024], Diet was also shown to yield interesting identifiable guanrantees [Reizinger et al., 2025], laying the theoretical foundation for RelGeo(Diet), where we employ the resulting geometry.
One can consider such a scenario [Reizinger et al., 2025]: some latent variables $\bm{z}$ are drawn from a vMF distribution, and pushed forward through a continuous and injective generator function $g$ to obtain the data $\bm{x}$. Remarkably, given only $\bm{x}$ without the knowledge of $g$, it is possible to (to some degree) recover the latent variables $\bm{z}$ through parameterizing a model and optimizing the instance discrimination loss as given in Equation 6. Specifically, suppose there is a finite set of vectors $\bm{v}_{c}$ on a unit sphere, each representing a class, and a finite set of instances. One instance belongs to exactly one class, and every class is employed by some instance. Additionally, the instance labels are chosen uniformly, and the latent variables $\bm{z}$ are drawn from a vMF distribution centered around the corresponding cluster vector $\bm{v}_{c}$ with concentration parameter $\kappa$.
Then, after the model is trained using the loss function as in Equation 6, when both $f$ and $w$ are not unit-normalized, $f\circ g$ is linear. This can be proven rigorously by expanding upon the theoretical framework of non-linear ICA [Hyvärinen et al., 2023]. As such, we propose to utilize $f\circ g$ to form the representations. For further technical details on the assumptions and additional results, we refer interested readers to Reizinger et al. [2025].
Assuming spherical geometry, the distance between two points $\bm{x}$ and $\bm{y}$ can be computed as
$d(\bm{x},\bm{y})=\arccos\left(\frac{\bm{x}^{\top}\bm{y}}{\left\lVert\bm{x}\right\rVert\left\lVert\bm{y}\right\rVert}\right).$
In the above formula, points that do not precisely lie on the unit sphere are effectively projected onto it. Interestingly this bears a strong resemblance to the cosine distances as used in the original paper on relative representations [Moschella et al., 2023].
A.2.2 Why it works
Prior work on representational alignment has shown that representations from different models can often be approximately aligned using simple transformations, e.g. linear, orthogonal or locally linear maps. Even when models are trained independently – with differrent architectures, modalities, or datasets that nonetheless share an underlying structure – they tend to learn similar representations, suggesting convergence towards a shared encoding of entities [Huh et al., 2024]. From a theoretical standpoint, identifiability results [Roeder et al., 2021] imply that if two discriminative models learn the same likelihood function, their internal representations must be equivalent up to a linear transformation. However, this ideal scenario rarely holds exactly in practice: training dynamics, nuisance factors, and unmodeled variability can all introduce distortion. In our case, it may be too strong to assume that two models learn different parameterizations of an identical manifold. Instead, we adopt a weaker assumption, that they do so up to some bounded distortion. Recent theoretical work has begun to explore relaxations of strict identifiability to account for such bounded distortions [Nielsen et al., 2025]. Integrating these relaxations into our Riemannian framework presents a promising direction for future work.
In general, the few theoretical results available, e.g. [Roeder et al., 2021], often rely on unrealistic assumptions, e.g. proofs in axiomatic settings, infinite-data regimes, or the requirement that two models learn exactly the same likelihood function. In our view, a meaningful first step toward bringing theory and practice is to relax these assumptions, as initiated in [Nielsen et al., 2025], and begin to model more realistic scenarios, e.g. including the dynamics introduced by model training.
We believe that Riemannian methods can play a key role in this direction to try to capture local alignments beyond linear global transformations of the space as considered in Roeder et al. [2021] and possibly accounting for distortions measured in the linear space in practice. Nevertheless, we remark that neural networks could find qualitatively different solutions [Pascanu et al., 2025] and that the union of manifolds hypothesis might be more appropriate for modeling image data [Brown et al., 2023].
A.3 Additional details
A.3.1 Mean Reciprocal Rank
@@
Mean Reciprocal Rank (MRR) is a commonly used metric to evaluate the performance of retrieval systems, and has been used to evaluate the capabilities of representations for instance discrimination [Moschella et al., 2023]. It measures the effectiveness of a system by calculating the rank of the first relevant item in the search results for each query.
@@
    $\text{MRR}=\frac{1}{Q}\sum_{i=1}^{Q}\frac{1}{r_{i}}.$
@@
Similar to stitching accuracies, MRR is generally asymmetric. However, it can also be made symmetric. Specifically, as MRR is calculated based on a distance matrix $D$, one can make the distance matrix symmetric by setting $D=\frac{1}{2}\left(D^{\top}+D\right)$. In Section 4.2.1 we reported the symmetric version. Otherwise we report both the original version and the symmetric version, and discriminate between these two by explicitly indicating it when it is symmetric.
A.3.2 Architectural details
We provide in Table 3 the architectural details of the convolutional autoencoders employed in experiments in Figures 3 and 4.
Table 3: Architecture of the convolutional autoencoders.
@@ A.3.2 Architectural details
A.3.3 RelGeo(Diet) augmentations
@@ A.3.4 Compute resources
The autoencoder stitching and retrieval experiments were conducted on a single NVIDIA RTX 3080TI GPU. Experiments involving vision foundation models were run on a compute cluster, each job using a single NVIDIA A100 GPU and 10 CPU cores, with runtimes of several hours. Preliminary experiments required additional resources, and in total we estimate having used several hundred GPU hours.
Further ablation studies on the running times can be found in Section A.4.10.
A.3.5 Geodesic approximation
@@
We trained a lightweight convolutional autoencoder (see Table 6) on both MNIST and CIFAR-10 to obtain the latent representations used in our experiments. For MNIST, the first convolutional layer was adjusted to accept a single input channel; for CIFAR-10 it used three channels. Each model was trained for 30 epochs using the Adam optimizer [Kingma and Ba, 2017] with a batch size of 64. We set the learning rate to 0.001, and fixed a random seed of 42 to ensure reproducibility.
Energy computation
After training, we selected 10 samples per class (100 total) in label order from each dataset and encoded them to produce their latent encodings. True geodesics are computed using Stochman library [Detlefsen et al., 2021], which has Apache-2.0 license, which wraps the decoder into a pullback manifold, intializes a parameterized spline path between codes, and then optimizes its parameters to minimize the Riemannian energy. Geodesic energies are computed as in Eq. 2. Pairwise energies are computed and visualized in Figures 2 and 7, demonstrating the close agreement between the two measures under identical encoding and discretization settings. In Fig. 2, latent dimensions for MNIST and CIFAR are 64 and 128 respectively, while in Fig. 7, latent dimension is 2 for both datasets.
@@
[…]
(a) MNIST
[…]
(b) CIFAR-10
Figure 7: Pairwise latent‐space energy matrices for (a) MNIST and (b) CIFAR-10, with latent dimensionality 2. In each subfigure, the left heatmap shows the straight-line energy proxy and the right shows the full Riemannian geodesic energies. The Spearman rank correlation between the two measures is 0.99 for MNIST and $\rho=1.00$ for CIFAR-10, demonstrating near-perfect agreement.
@@
Figure 8: Impact of varying discretization levels on similarity and energy metrics for (a) MNIST and (b) CIFAR‑10 datasets. Each subplot shows how Spearman’s $\rho$, Pearson’s $r$, and Euclidean distance change as the number of discretization levels increases.
A.3.6 Autoencoder stitching and retrieval
@@
We trained the lightweight convolutional autoencoder (see Table 6) on MNIST, CIFAR-10, FashionMNIST with 5 different seeds, to obtain the latent representations used in our experiments. For MNIST and FashionMNIST the first convolutional layer was adjusted to accept a single input channel; for CIFAR-10 it used three channels. Each model was trained for 50 epochs, reaching convergence, using the Adam optimizer [Kingma and Ba, 2017] with a batch size of 64. We set the learning rate to 0.001.
A.3.7 Vision foundation models
@@
For CIFAR-100, we use the coarse labels. For SVHN, the objective is to predict the cropped digits. For CUB dataset, we use the version available at https://huggingface.co/datasets/birder-project/CUB_200_2011-WDS. Given the relatively small training set, we select $2000$ points as the validation set. When reporting aggregated MRR metrics in the tables, we always exclude the diagonal entries as these are generally (close to) $1$. For ImageNet-1k, we use the validation set and split it into the final train, val and test sets. Further details can be found in the provided code.
For all cases where we need to train classification heads, apart from the ones with Diet the heads are trained for $10$ epochs, while the ones with Diet are trained for $50$ epochs. The heads used to obtain the gometric information are trained using learning rate 5e45e-4 and batch size $64$, while the heads used for stitching was trained using learning rate 1e41e-4 and batch size $32$. We always use the Adam optimizer [Kingma and Ba, 2017].
@@
A.4 Additional results on Vision Foundation models
We provide additional results on vision foundation models. For ablation studies, we focus on the performances of the models on CUB dataset. We refer to accuracy as Accuracy, symmetricized MRR based on cosine as MRR Cosine Sym, symmetricized MRR based on cdist as MRR CDist Sym, MRR based on cosine as MRR Cosine and MRR based on cdist as MRR CDist.
A.4.1 Full results
We provide the heatmaps on the different datasets in Figure 9, Figure 10, Figure 11, Figure 12 and Figure 13.
Figure 9: Results on CIFAR-10. From top to bottom: Accuracy, MRR Cosine Sym, MRR CDist Sym, MRR Cosine, MRR CDist
Figure 10: Results on CIFAR-100. From top to bottom: Accuracy, MRR Cosine Sym, MRR CDist Sym, MRR Cosine, MRR CDist
Figure 11: Results on ImageNet-1k. From top to bottom: Accuracy, MRR Cosine Sym, MRR CDist Sym, MRR Cosine, MRR CDist
Figure 12: Results on CUB. From top to bottom: Accuracy, MRR Cosine Sym, MRR CDist Sym, MRR Cosine, MRR CDist
Figure 13: Results on SVHN. From top to bottom: Accuracy, MRR Cosine Sym, MRR CDist Sym, MRR Cosine, MRR CDist
A.4.2 Other evaluation metrics